# Seeing the Image: Prioritizing Visual Correlation by Contrastive Alignment

### v1

@@ Abstract
Existing image-text modality alignment in Vision Language Models (VLMs) treats each text token equally in an autoregressive manner. Despite being simple and effective, this method results in sub-optimal cross-modal alignment by over-emphasizing the text tokens that are less correlated with or even contradictory with the input images. In this paper, we advocate for assigning distinct contributions for each text token based on its visual correlation. Specifically, we present by contrasting image inputs, the difference in prediction logits on each text token provides strong guidance of visual correlation. We therefore introduce Contrastive ALignment (CAL), a simple yet effective re-weighting strategy that prioritizes training visually correlated tokens. Our experimental results demonstrate that CAL consistently improves different types of VLMs across different resolutions and model sizes on various benchmark datasets. Importantly, our method incurs minimal additional computational overhead, rendering it highly efficient compared to alternative data scaling strategies. Codes are available at github.com/foundation-multimodal-models/CAL.
@@ 1 Introduction
Despite these advancements, whether the current alignment strategy on existing image-text datasets performs satisfactorily is often less studied. The existing alignment strategy simply treats all text tokens equally in an auto-regressive manner. Although such a method has been proven to be simple and effective, many text tokens in the training labels are irrelevant to the visual inputs, which contribute little to image-text modality alignment. Figure 1(a) presents a sample drawn from the ShareGPT4V [16] dataset, where a large proportion of text tokens including unique, context presents little visual correlation. Treating these text tokens in equal weights results in ineffective training and can introduce negative effects by prioritizing more on fitting the distribution of these visually irrelevant tokens, rather than the image-text modality alignment.
@@
Moreover, there also exists a proportion of tokens that contradicts visual conditions, which is inevitable in model generated datasets [7, 16, 13, 15, 14], also presented in Figure 1(a). In particular, we conduct human evaluations on the broadly used GPT-assisted datasets including ShareGPT4V and the detail caption subset of LLaVA-Instruct in Figure 1(b) by sampling 100 samples from each dataset. We score each sample by 0 and 1 based on whether there exist text tokens that contradict with visual input, and the score is averaged by three annotators. We found approximately half of the sampled datasets contain visually contradictory tokens in both datasets. Imitating the text distribution on these contradictory tokens further harms the image-text modality alignment. Consequently, recent evaluations on existing VLMs [33, 34, 5, 35, 6, 8] present the shortcomings of current alignment strategy from various aspects, including hallucination [36, 37, 33] and responding without depending on visual conditions [27, 38].
@@
Fortunately, inspired by recent training-free visual contrastive decoding researches [36, 33, 39], we present that the visual correlation can be directly indicated by contrasting input image conditions. In particular, we investigate the change in prediction logits of text tokens with or without the image input and observe strong relevance between the logit change of each text token and its visual correlation. We therefore propose Contrastive ALignment (CAL), which is a surprisingly simple re-weighting strategy to prioritize the training of text tokens that are highly correlated with the input image to enhance image-text modality alignment. Experiments have shown that our proposed method can improve leading VLMs of different kinds including LLaVA-1.5/LLaVA-NeXT [7, 6, 10], MiniGemini(MGM)/MGM-HD [11], across different resolution and model size on various types of benchmarks including visual question answering, captioning and grounding. Especially, CAL on LLaVA-Next-13B [10] can bring an impressive performance of 1.7 ANLS on VQA [26], 3.4 relaxed accuracy on VQA [25], 2.2 CIDEr [40] on COCO [30] and 6.3 CIDEr on TextCaps [29], 0.6/0.7 IoU on validation/test set of RefCOCOg [32]. Moreover, our method introduces little computational overhead, with only one auxiliary forward operation in each training step. The lightweight feature while the impressive performance of CAL brings current VLMs to a new stage, highlighting the importance of a delicate image-text modality alignment strategy design. We further conduct extensive qualitative analysis for CAL and present the improved ability of OCR recognition and image-captioning.
@@ 3.1 Experimental Setup
In this paper, we verify our proposed CAL on two leading model structures: LLaVA-1.5/LLaVA-NeXT [6, 10] and Mini-Gemini/Mini-Gemini-HD [11]. LLaVA-1.5 uses CLIP-pretrained ViT-L as the visual encoder. For resolution scaling, LLaVA-NeXT employs a simple while adaptive image cropping strategy, encodes each image and concatenates them in one single sequence. Mini-Gemini (MGM) further introduces a LAION-pretrained ConvNeXt-L [22, 41] for high-resolution refinement. For MGM/MGM-HD/LLaVA-1.5, we follow the same setting as the original paper as it is public available, where the learning rate for the PT stage is set to and the IT stage is set to for both Vicuna-7B and Vicuna-13B. For LLaVA-NeXT, where only the evaluation code is made public, we reproduce LLaVA-NeXT with the same learning rate as MGM, and set the learning rate of ViT to 1/10 of the base learning rate (our reproduction presents on-par performance with the original paper/blog. We present a comparison of our reproduction results with those of the original papers in Appendix A.1). We also set the lower bound and upper bound in Equation 3 to 1 and 5 respectively, and we set in Equation 4 to 3 for all experiments. We use 16 A100 for experiments, except for 8 GPUs in LLavA-1.5/Gemma-2B and 32 GPUs in MGM-HD-13B.
@@ 3.3 Ablation Studies
Stages of Conducting CAL within VLM Training. CAL can be integrated into both the PreTraining (PT) stage and the Instruction Tuning (IT) stage in existing VLMs. In this section, we investigate which stage benefits the most from CAL in Table 4. Integrating CAL into the IT stage contributes most of the improvement in all listed benchmarks, and CAL in the PT stage further enhances the performance with notable improvement on MMT-Bench and OCR-Bench.
@@ 4 Related Work
LLMs [1, 52, 2, 53, 4] have made significant strides in Natural Language Processing (NLP) tasks, including text generation and question-answering, paving the way for VLMs that integrate vision ability with LLMs. In the realm of visual language learning, CLIP [17] has set a milestone by employing extensive image-text pair contrastive learning to achieve multimodal alignment. Recently, numerous VLMs [5, 7, 8, 54, 34, 55, 56, 57, 58, 35, 59, 60] have leveraged the robust capabilities of LLMs for cross-modal understanding and generation tasks. Models like BLIP-2 [5] and MiniGPT-4 [8] have improved cross-modal alignment through comprehensive image-text pair pre-training. LLaVA [7] has further advanced its comprehension of complex prompts via refined instruction fine-tuning. Additionally, recent research [9, 10, 11] has incorporated higher resolution input images and longer sequences to enhance VLMs’ understanding capabilities. Mini-Gemini (MGM) [11] introduces a LAION-pretrained ConvNeXt-L [22] for high-resolution refinement.
@@
Image-text modality alignment has long been regarded as the core problem in cross-modal understanding and generation tasks. Traditional image-text alignment strategies include both contrastive learning across different modalities and generative learning that train text tokens in an autoregressive manner [17, 61, 62]. The combination of both techniques is also proven to be effective in the early era of VLMs, where BLIP [5] proposes a multi-stage alignment strategy, with contrastive learning in early alignment and generative learning in the latter stage. However, in recent researches [7, 10], contrastive learning is discarded for being redundant in image-text modality alignment of VLMs, and researchers propose to enhance the cross-modal alignment through dataset scaling and image resolution scaling [53, 9, 63, 10]. Despite being simple in application, the existing generative alignment method simply treats each text token with equal importance, resulting in sub-optimal alignment performance.
@@ 5 Limitation
The selection of lower bounds and upper bounds in Equation 3 are empirically decided based on the frequency of the prediction logits, which could be extended to more adaptive settings in further explorations. Nevertheless, the simple while broadly effective nature of CAL indicates the importance of a delicate image-text modality alignment strategy for leading VLM structures.
@@ A.3 Ablations for Image Contrasting Conditions
A.4 Computational Overhead.
@@
Each iteration of our proposed method requires forwarding text tokens twice to effectively enhance image-text modality alignment. Table 10 presents the training time for the instruct-tuning stage on LLaVA-NeXT, with each experiment conducted on 16 A100 GPUs. CAL introduces approximately 20% computational overhead on both LLaVA-NeXT-7B and LLaVA-NeXT-13B. Our implementation currently uses an attention mask to neutralize the effect of the image tokens, meaning that the image tokens are still forwarded twice. This additional burden could be significantly reduced by completely removing the image tokens.
@@ B.2 Image-text Modality Alignment Visualization
![[Uncaptioned image]](extracted/5625265/Styles/figures/cases/ocr_case3.png)
![[Uncaptioned image]](extracted/5625265/Styles/figures/cases/ocr_case4.png)
![[Uncaptioned image]](extracted/5625265/Styles/figures/cases/short_cap_case1.png)
![[Uncaptioned image]](extracted/5625265/Styles/figures/cases/short_cap_case2.png)
![[Uncaptioned image]](extracted/5625265/Styles/figures/cases/short_cap_case6.jpg)
![[Uncaptioned image]](extracted/5625265/Styles/figures/cases/short_cap_case7.jpg)

### v2

@@ Abstract
Existing image-text modality alignment in Vision Language Models (VLMs) treats each text token equally in an autoregressive manner. Despite being simple and effective, this method results in sub-optimal cross-modal alignment by over-emphasizing the text tokens that are less correlated with or even contradictory with the input images. In this paper, we advocate for assigning distinct contributions for each text token based on its visual correlation. Specifically, we present by contrasting image inputs, the difference in prediction logits on each text token provides strong guidance of visual correlation. We therefore introduce Contrastive ALignment (CAL), a simple yet effective re-weighting strategy that prioritizes visually correlated tokens. Our experimental results demonstrate that CAL consistently improves different types of VLMs across different resolutions and model sizes on various benchmarks. Importantly, our method incurs minimal additional computational overhead, rendering it highly efficient compared to alternative data scaling strategies.
@@ 1 Introduction
Despite these advancements, whether the current alignment strategy on existing image-text datasets performs satisfactorily is often less studied. Existing alignment strategies simply treat all text tokens equally in an auto-regressive manner. Although such a method has been proven to be simple and effective, many text tokens exhibit limited relevance to the visual inputs, which contribute little to image-text modality alignment. Figure 1(a) presents a sample drawn from the ShareGPT4V [16] dataset, where a large proportion of text tokens including unique, context presents little visual correlation. Treating these text tokens in equal weights results in ineffective training and can introduce negative effects by prioritizing more on fitting the distribution of these visually irrelevant tokens, rather than the image-text modality alignment.
@@
Moreover, there also exists a proportion of tokens that contradicts visual conditions, which is inevitable in model generated datasets [7, 16, 13, 15, 14], presented in Figure 1(a). In particular, we conduct human evaluations on the broadly used GPT-assisted datasets including ShareGPT4V and the detail caption subset of LLaVA-Instruct in Figure 1(b) by sampling 100 samples from each dataset. We score each sample by 0 and 1 based on whether there exist text tokens that contradict with visual input, and the score is averaged by three annotators. We found approximately half of the sampled datasets contain visually contradictory tokens in both datasets. Imitating the text distribution on these contradictory tokens further harms the image-text modality alignment. Consequently, recent evaluations on existing VLMs [33, 34, 5, 35, 6, 8] present the shortcomings of current alignment strategy from various aspects, including hallucination [36, 37, 33] and responding without depending on visual conditions [27, 38].
@@
Fortunately, inspired by recent training-free visual contrastive decoding researches [36, 33, 39], we present that the visual correlation can be directly indicated by contrasting input image conditions. In particular, we investigate the change in prediction logits of text tokens with or without the image input and observe strong relevance between the logit change of each text token and its visual correlation. We therefore propose Contrastive ALignment (CAL), which is a surprisingly simple re-weighting strategy to prioritize the training of text tokens that are highly correlated with the input image to enhance image-text modality alignment. Experiments have shown that our proposed method can improve leading VLMs of different kinds including LLaVA-1.5/LLaVA-NeXT [7, 6, 10], MiniGemini(MGM)/MGM-HD [11], across different resolution and model size on various types of benchmarks including visual question answering, captioning and grounding. Especially, CAL on LLaVA-Next-13B [10] can bring an impressive performance of 1.7 ANLS on VQA [26], 3.4 relaxed accuracy on VQA [25], 2.2 CIDEr [40] on COCO [30] and 6.3 CIDEr on TextCaps [29], 0.6/0.7 IoU on validation/test set of RefCOCOg [32]. Moreover, our method introduces little computational overhead, with one auxiliary gradient-free forward operation in each training step. The lightweight feature while the impressive performance of CAL brings current VLMs to a new stage, highlighting the importance of a delicate image-text modality alignment strategy design. We further conduct extensive qualitative analysis for CAL and present the improved ability of OCR recognition and image-captioning.
@@ 3.1 Experimental Setup
In this paper, we verify our proposed CAL on two leading model structures: LLaVA-1.5/LLaVA-NeXT [6, 10] and Mini-Gemini/Mini-Gemini-HD [11]. LLaVA-1.5 uses CLIP-pretrained ViT-L as the visual encoder. For resolution scaling, LLaVA-NeXT employs a simple while adaptive image cropping strategy, encodes each image and concatenates them in one single sequence. Mini-Gemini (MGM) further introduces a LAION-pretrained ConvNeXt-L [22, 41] for high-resolution refinement. For MGM/MGM-HD/LLaVA-1.5, we follow the same setting as the original paper as it is public available, where the learning rate for the PT stage is set to and the IT stage is set to for both Vicuna-7B and Vicuna-13B. For LLaVA-NeXT, where only the evaluation code is made public, we reproduce LLaVA-NeXT with the same learning rate as MGM, and set the learning rate of ViT to 1/10 of the base learning rate (our reproduction presents on-par performance with the original paper/blog. We present a comparison of our reproduction results with those of the original papers in Appendix A.1). We also set the lower bound and upper bound in Equation (3) to 1 and 5 respectively, and we set in Equation (4) to 3 for all experiments. We use 16 A100 for experiments, except for 8 GPUs in LLavA-1.5/Gemma-2B and 32 GPUs in MGM-HD-13B.
@@ 3.3 Ablation Studies
Stages of Conducting CAL within VLM Training CAL can be integrated into both the PreTraining (PT) stage and the Instruction Tuning (IT) stage in existing VLMs. In this section, we investigate which stage benefits the most from CAL in Table 4. Integrating CAL into the IT stage contributes most of the improvement in all listed benchmarks, and CAL in the PT stage further enhances the performance with notable improvement on MMT-Bench and OCR-Bench.
@@ 4 Related Work
LLMs [1, 52, 2, 53, 4] have made significant strides in Natural Language Processing (NLP) tasks, including text generation and question-answering, paving the way for VLMs that integrate vision ability with LLMs. In the realm of visual language learning, CLIP [17, 54] has set a milestone by employing extensive image-text pair contrastive learning to achieve multimodal alignment. Recently, numerous VLMs [5, 7, 8, 55, 34, 56, 57, 58, 59, 35, 60, 61] have leveraged the robust capabilities of LLMs for cross-modal understanding and generation tasks. Models like BLIP-2 [5] and MiniGPT-4 [8] have improved cross-modal alignment through comprehensive image-text pair pre-training. LLaVA [7] has further advanced its comprehension of complex prompts via refined instruction fine-tuning. Additionally, recent research [9, 10, 11] has incorporated higher resolution input images and longer sequences to enhance VLMs’ understanding capabilities. Mini-Gemini (MGM) [11] introduces a LAION-pretrained ConvNeXt-L [22] for high-resolution refinement.
@@
Image-text modality alignment has long been regarded as the core problem in cross-modal understanding and generation tasks. Traditional image-text alignment strategies include both contrastive learning across different modalities and generative learning that train text tokens in an autoregressive manner [17, 62, 63, 64]. The combination of both techniques is also proven to be effective in the early era of VLMs, where BLIP [5] proposes a multi-stage alignment strategy, with contrastive learning in early alignment and generative learning in the latter stage. However, in recent researches [7, 10], contrastive learning is discarded for being redundant in image-text modality alignment of VLMs, and researchers propose to enhance the cross-modal alignment through dataset scaling and image resolution scaling [53, 9, 65, 10]. Despite being simple in application, the existing generative alignment method simply treats each text token with equal importance, resulting in sub-optimal alignment performance.
@@ 5 Limitation
The selection of lower bounds and upper bounds in Equation (3) are empirically decided based on the frequency of the prediction logits, which could be extended to more adaptive settings in further explorations. Nevertheless, the simple while broadly effective nature of CAL indicates the importance of a delicate image-text modality alignment strategy for leading VLM structures.
@@ A.3 Ablations for Image Contrasting Conditions
A.4 Ablations for Average Pooling
| ChartVQA | DocVQA | SQAI | COCO Caption | TextCaps | OCRB. | Refcocogval | |
| w/ AvgPool | 67.2 | 80.1 | 71.5 | 120.6 | 124.4 | 574 | 80.4 |
| w/o AvgPool | 66.3 | 79.5 | 72.4 | 116.7 | 123.8 | 581 | 79.5 |
A.5 Ablations for Pre-trained Model
| Model | Pretrain | ChartVQA | DocVQA | SQAI | COCO Caption | TextCaps | OCRB. | Refcocogval |
| Baseline | Original | 63.8 | 78.4 | 71.8 | 118.5 | 118.2 | 553 | 79.8 |
| CAL | Original | 67.2 | 80.1 | 71.5 | 120.6 | 124.4 | 574 | 80.4 |
| CAL | Baseline | 66.3 | 79.5 | 72.4 | 116.7 | 123.8 | 581 | 80.2 |
One assumption of CAL is that after training a simple projector only, the VLM is capable of distinguishing visually correlated tokens. In the first phase of VLM training, it is common practice to freeze the ViT and the LLM, and only train a projector to align their features. In the second phase, we finetune the model using high-quality data to enable it to answer image-related questions. To validate this assumption, we finetune CAL models using two types of pre-trained versions: one with the original pre-trained model (only train the projector) and one with the fully trained baseline model (after the finetuning phase). As shown in Table 11, we compare the performance of these two versions and found no significant differences in performance.
A.6 Computational Overhead.
@@ A.5 Ablations for Pre-trained Model
Each iteration of our proposed method requires forwarding text tokens twice to effectively enhance image-text modality alignment. Table 12 presents the training time for the instruct-tuning stage on LLaVA-NeXT, with each experiment conducted on 16 A100 GPUs. CAL introduces approximately 20% computational overhead on both LLaVA-NeXT-7B and LLaVA-NeXT-13B. Our implementation currently uses an attention mask to neutralize the effect of the image tokens, meaning that the image tokens are still forwarded twice. This additional burden could be further reduced by completely removing image tokens.
@@ B.2 Image-text Modality Alignment Visualization
![[Uncaptioned image]](extracted/5978023/Styles/figures/cases/ocr_case3.png)
![[Uncaptioned image]](extracted/5978023/Styles/figures/cases/ocr_case4.png)
![[Uncaptioned image]](extracted/5978023/Styles/figures/cases/short_cap_case1.png)
![[Uncaptioned image]](extracted/5978023/Styles/figures/cases/short_cap_case2.png)
![[Uncaptioned image]](extracted/5978023/Styles/figures/cases/short_cap_case6.jpg)
![[Uncaptioned image]](extracted/5978023/Styles/figures/cases/short_cap_case7.jpg)